OpenStudy (anonymous):

Simplify: ( y^−7 × y^−3 )^−1

OpenStudy (anonymous):

\[(y{}^{\wedge}-7\times y{}^{\wedge}-3){}^{\wedge}-1 \]\[((y{}^{\wedge}-7) (y{}^{\wedge}-3)){}^{\wedge}-1 \]\[\frac{1}{(y{}^{\wedge}-7) (y{}^{\wedge}-3)} \]\[\frac{1}{\left(\frac{1}{y^7}\right) \left(\frac{1}{y^3}\right)} \]\[\frac{1}{\frac{1}{y^{10}}} \]\[y^{10} \]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You are welcome.